OpenStudy (anonymous):

simplify each product. 1.(s+7t^2)^2 2.p^5-8q^3)^2 3.e^4+f^2)^2

OpenStudy (diyadiya):

\[(s+7t^2)^2\] Use \[(a+b)^2=a^2+2ab+b^2\] \[Here, \ a=s \ \ and \ \ b= 7t^2\]